# Hybrid Swin Attention Networks for Simultaneously Low-Dose PET and CT Denoising

### v4

@@ 3.2.2 Quantitative evaluation results on LDPET dataset
We have used the max-min normalization method to normalize images for training. An average test results on all 9 patients are shown in Table 4. As shown in the table, our method outperforms the other approaches and achieves a PSNR of 32.73. Although its performance is 0.31 dB lower than Unet, it achieves a comparable SSIM with substantially fewer parameters. Figure LABEL:LDPET_violin shows the violin plots for the 9 patients in the test set. Our model demonstrates more stable PSNR and RMSE performance across patients compared with Unet. Note that in the PET denoising experiment, we did not use the Sobel loss due to convergence issues.
Figure LABEL:LDPET_fig shows the denoised results of an abdominal image. As observed, CTformer and Swin-Unet produce grid-like artifacts in the image. RED-CNN and SwinIR fail to capture the image patterns. U-Net and HSANet appear to capture them better than the other methods, although their results are overly smooth.
@@
ours 32.73 0.96 6.9 5.4G 0.61M
@@ 4 Discussion and conclusion
To address the challenges posed by complex, structured noise that varies across patients, scanners, and anatomical regions, and to maintain a delicate balance between over-smoothing and under-denoising, which can leave residual artifacts, we developed HASNet, which integrates our proposed efficient attention mechanisms with a novel upsampling strategy, the HIC module. This design leads to state-of-the-art denoising performance, achieving the highest PSNR and lowest RMSE among all compared methods. Beyond accuracy, HASNet demonstrates remarkable stability across multiple runs, with minimal performance variation. Additionally, it is highly efficient: HASNet contains significantly fewer parameters than conventional models and requires substantially less GPU memory, making it not only accurate and robust but also lightweight and practical for real-world clinical deployment.
@@ Acknowledgments
We are grateful for access to the University of Heidelberg’s IWR HPC service for running simulations used in this study.
Appendix A Example Appendix Section

### v5

@@ 3.2.2 Quantitative evaluation results on LDPET dataset
We have used the max-min normalization method to normalize images for training. An average test results on all 9 patients are shown in Table 4. As shown in the table, our base model outperforms the other approaches and achieves a PSNR of 32.73, except Unet. Although its performance is 0.31 dB lower than Unet, it achieves a comparable SSIM with substantially fewer parameters. The model was further scaled up by increasing the number of residual convolutional layers with skip-connections and incorporating two residual Swin-Transformer bottleneck blocks, each containing four bottleneck units. In addition, since the ESGA module within the Swin-Transformer block did not improve performance, the original Swin-Transformer module was retained. Figure 8(a) shows the violin plots of the scaled large HSANet for the 9 patients in the test set. Our model demonstrates slightly better results on PSNR and RMSE performance across patients compared with Unet. Note that in the PET denoising experiment, we did not use the Sobel loss due to convergence issues.
Figure 8(b) shows the denoised results of an abdominal image. As observed, CTformer and Swin-Unet produce grid-like artifacts in the image. RED-CNN and SwinIR fail to capture the image patterns. U-Net and HSANet appear to capture them better than the other methods, although their results are overly smooth.
@@
HSANet(base) 32.73 0.96 6.9 5.4G 0.61M
HSANet(large) 33.11 0.96 6.55 10.4G 3.88M
(a) Quantitative (a)PSNR, (b)SSIM and (c)RMSE of 9 different PET patients on test set. Red points are average. Width of violin plot represent the density of data at each value. HSANet(b) and HSANet(l) represent base and large model respectively. Quartiles are shown as thick lines inside the violin plot
(b) Results of abdomen image for comparison. (a)LDPET, (b)RED-CNN,(c)Swin-Unet, (d)SwinIR, (e)CTformer, (f)Unet, (g)HSANet large, (h)FDPET
3.3 Scalability of Swin-Unet
In Swin-Unet, the authors showed that increasing Swin-Transformer block numbers (network depth) does not significantly improve segmentation performance but considerably increases computational cost [cao2022swin]. In line with Xiao et al. [xiao2021early], who reported that the limited locality bias of Swin-Transformers makes them suboptimal for early feature extraction, we scaled up only the bottleneck stage using residual Swin-Transformer blocks while keeping the shallow Swin-Transformer layers unchanged, which led to improved performance. Furthermore, the number of convolutional blocks was increased from two to four in both the encoder and decoder, with skip connections added to improve feature reuse. The comparison detail shows in Table 5. Note that decay factor $\gamma$ here is set to 0.85 for scaled HSANet model.
Table 5: scalability of HSANet with residual Swin-Transformer bottleneck
Swin-Transformer bottleneck
PSNR 33.11 32.88
@@ 4 Discussion and conclusion
To address the challenges posed by complex, structured noise that varies across patients, scanners, and anatomical regions, and to maintain a delicate balance between over-smoothing and under-denoising, which can leave residual artifacts, we developed HASNet, which integrates our proposed efficient attention mechanisms with a novel upsampling strategy, the HIC module. This design leads to state-of-the-art denoising performance, achieving the highest PSNR and lowest RMSE among all compared methods. Beyond accuracy, HASNet demonstrates remarkable stability across multiple runs, with minimal performance variation. Additionally, it is highly efficient, particularly on the CT dataset. It contains significantly fewer parameters and requires GPU memory comparable to conventional models, making it not only accurate and robust but also lightweight and practical for real-world clinical deployment.
@@ Acknowledgments
Data used in the preparation of this article were obtained from the University of Bern, Department of Nuclear Medicine, and the School of Medicine, Ruijin Hospital. The investigators at these institutions contributed to the design and implementation of the data collection and/or provided the data but did not participate in the analysis or writing of this manuscript.
Data availibility
The datasets are public available online. All deep learning methods were implemented using Pytorch (https://pytorch.org/). The custom script for this study will be available at https://github.com/Christian-lyc/HSANet.
Appendix A Appendix figures